# Proposal for an experiment to verify Wigner’s rotation at non-relativistic speeds with massive spin-$1/2$ particles

### v3

@@ V Acknowledgments
We would like to thank the anonymous reviewers for raising a number of important points, as well as for helpful suggestions that improved the quality of the paper. VP acknowledges partial support by the Estonian Research Council (Eesti Teadusagentuur, ETAG) through grants PSG489 and PRG946. CP acknowledges support by the excellence cluster QuantumFrontiers of the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under Germany’s Excellence Strategy, EXC-2123 QuantumFrontiers, 390837967; support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project Number 420243324 and by the Transilvania Fellowships for Visiting Professors grant 2024 of the Transilvania University of Brasov.

### v4

@@ V Acknowledgments
We would like to thank the anonymous reviewers for raising a number of important points, as well as for helpful suggestions that improved the quality of the paper. VP acknowledges partial support by the Estonian Research Council (Eesti Teadusagentuur, ETAG) through grants PSG489 and PRG946. CP acknowledges support by the excellence cluster QuantumFrontiers of the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under Germany’s Excellence Strategy, EXC-2123 QuantumFrontiers, 390837967; support by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG), Project Number 420243324 and by the Transilvania Fellowships for Visiting Professors grant 2024 of the Transilvania University of Brasov.